(D2-2D+3)y=x2-1


1
Expert's answer
2021-06-23T14:13:39-0400

"(D^2 - 2D +3) y = x^2 - 1"

The differential equation can be written as

"y'' - 2y' +3y = x^2 -1"

The auxiliary equation of the homogenous part is

"m^2-2m+3 = 0"

Solving the quadratic equation, we have that

"m = 1 \pm \sqrt{2}i"

Hence the solution of the homogenous part is

"y = e^x(c_1\cos \sqrt{2}x + i \,c_2\sin \sqrt{2}x)"

We use the method of undetermined coefficient to solve the other part.

Suppose

"y = ax^2 + bx + c"

Then

"y' = 2ax + b"

"y'' = 2a"

Substituting into the given differential equation, we have

"2a-2(2ax+b)+3(ax^2+bx+c) = x^2 - 1"

Comparing the coefficients of "x^2" we have

"3a = 1 \implies a = \dfrac{1}{3}"

Comparing the coefficients of x, we have

"-4a +3b = 0 \implies b = \dfrac{4}{9}"

Comparing the constants, we have"2a -2b+3c = -1 \implies c = - \dfrac{7}{27}"

Hence, we have that the solution for this part is "y= \frac{1}{3}x^2 + \frac{4}{9}x - \frac{7}{27}"

So the general solution is

"y = e^x(c_1\cos \sqrt{2}x + i \,c_2\sin \sqrt{2}x) +\frac{1}{3}x^2 + \frac{4}{9}x - \frac{7}{27}"